OpenStudy (anonymous):

algebra 1/II question!! I need to factor this: 4y2-5y-4

OpenStudy (crashonce):

you cant

OpenStudy (crashonce):

use quadratic formula

OpenStudy (crashonce):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

\[x=-b \pm \sqrt{b^2-4ac}/2a\]

OpenStudy (bradely):

4y^2-5y-4 Use quadratic formula y = (-(-5)±sqrt((-5)^2-4(4)(-4)))/2(4) =(5±sqrt(25+64)/8 =(5±sqrt(89)/8 Source: http://www.mathskey.com/question2answer/